Two point charges q and 4q are held at separation .At what distance the electric field due to them is zero?

Grade:12

2 Answers

GOROJANAPU KIRAN
10 years ago

q-----x---,---(d-x)------4q

x^2/(d-x)^2=1/4

x/(d-x)=1/2

d/x-1=2

d/x=3/2

x=2/3d

vamsi
10 years ago
q-----x---,---(d-x)------4q x^2/(d-x)^2=1/4 x/(d-x)=1/2 d/x-1=2 d/x=3/2 x=2/3d